# Superconducting spintronics with electron symmetry filtering and interfacial spin-orbit coupling

### v1

@@ IV.2.1 Superconductivity-induced change in the IP magnetic anisotropy
The most effective way of exploring the in-plane MCA landscape of the S/F/F junctions were magnetic field rotations, in which we used a field that was greater than the soft F layer coercive field but lower than the hard F layer one (see Figure 3b). The range of 70 to 120 Oe worked best in our case, depending on the sample. Using this range of fields ensured that the hard F layer remained fixed while the soft one could freely rotate its magnetization following the applied field direction, and revealing its own MCA in the process. This is shown in Figure 3b, where we can distinguish the two easy axes aligned with the [100] and [010] crystallographic directions as the soft F layer moves, setting the junction in the parallel (P), perpendicular-in-plane (PIP) and anti-parallel (AP) magnetic configurations.
@@
We used the experimental resistances to approximate the exact angle between both F layers in any situation where we could find different resistance values (found both above and under the critical temperature of vanadium). We used these resistances, for known magnetic configurations, to calibrate a modified version of the Slonczewski model Slonczewski, which can accurately model the S/F/F junctions due to their large effective spin polarization Martinez2018; Parkin2004; Yuasa2004.
@@
Figure 8: Resistance vs IP magnetic field rotation angle for a V/MgO/Fe/MgO/Fe/Co junction, for different temperatures from above to below $T_{C}$, when the rotation is started from the P (a-c) and the AP (d-f) magnetic configurations, with the corresponding crystalline axis indicated. The inset helps visualizing the experiment, defining the angles between the two F layers ($\Phi_{\text{F}}$) and the one of the applied field ($\Phi_{\text{H}}$). (b) and (e) show 10 histograms from the data in (a) and (d) respectively. (c) and (f) take the data from (b) and (e) during the second half of the magnetic field rotation, representing each histogram’s count vs temperature around the intermediate (normal state hard axis) states: $\text{AP}+45^{\circ}$ or the $\overline{1}10]$ axis, $\text{AP}+90^{\circ}=\text{PIP}$ or [010], and $\text{AP}+135^{\circ}$ or [110]. Figure adapted from ref. GonzalezRuano2020 with the necessary permissions.
@@
The magnetization rotations were then performed at different temperatures, from above to below $T_{C}$, in several samples. In Figure 8 we show above gap conductance versus applied magnetic field angle on two different samples, with the correspondent relative magnetization angle analysis. While the resulting TMR is not affected by temperature during the first half of the rotation, the second half has noticeable changes of the relative magnetization angle below $T_{C}$. In particular, we see new free layer magnetization configurations emerge in the superconducting state. According to the calibrated model, it seems to be oriented near the hard axes of the soft F layer, 45 from the easy directions. These new configurations disappear when increasing temperature and approaching $T_{C}$.
@@
Before continuing to fit the experimental results to the theoretical model, we want to discard some extrinsic effects. In the IP situation, the main candidate is the formation of domain walls (DWs). To asses their possible role in changing the resistance of the junctions during the field rotations (which could potentially invalidate our macrospin approach with the Slonczewsky model), we used the MuMax3 mumax micromagnetic simulation software to compare the formation of DWs in the [100] and [110] directions. The results showed that these DWs, in case they are present, should preferentially pin the magnetization along the [100] direction, and form less in the [110] one (see Ref. gonzalez-prb-20). Since we observe a high spin polarization when the magnetization is saturated along the [100] axis, which indicates that no DWs are present, we can be reasonably confident that no DWs should appear during the rest of the rotations.
@@
Once extrinsic factors were taken care of, we turn to the theoretical model Johnsen2019, adapting it to our system. As mentioned, the main idea is that the invariance of the superconducting proximity effect to the angle of the F layer magnetization is broken in the presence of SOC Jacobsen2015; Johnsen2019. Formally, the model is based on a tight-binding Bogolioubov-de Gennes Hamiltonian, which includes terms for electron hopping both within and between the different layers, interfacial Rashba SOC, spin exchange splitting for the Fe layers, and conventional superconductivity in vanadium. The total free energy is calculated by taking into account the magnetization-dependent superconducting proximity effect, bi-axial in-plane MCA, and a weak anti-ferromagnetic coupling between the F layers that was observed experimentally. The result is shown in Figure 9, where the free energy is plotted vs the angle of the magnetization ($\Phi_{\text{F}}$) under different temperatures. This nicely reproduces the new 45 easy axis below $T_{C}$. For more calculation details, seeGonzalezRuano2020.
@@
Figure 9: (a) Free energy $F$ vs in-plane magnetization angle $\Phi_{\text{F}}$ for different temperatures around $T_{C}$. (b) Sketch of the physical mechanism behind the S-induced change in MCA: Above $T_{C}$, V is a normal metal, while the Fe layer has a 4-fold in-plane MCA (yellow line). Below $T_{C}$, V is superconducting and influences the Fe with a magnetization-dependent penetration of Cooper pairs. The generation of triplets is lower for the $\overline{1}\overline{1}0]$ direction, maximizing the superconducting condensate energy gain. This in turn modifies the effective MCA of the Fe layer (cyan line), allowing the magnetization to reorient towards the $\overline{1}\overline{1}0]$ axis (blue arrow). Figure adapted from ref. GonzalezRuano2020 with the necessary permissions.
@@ IV.2.2 Superconductivity-assisted change of the OOP magnetic anisotropy
We turn now to describe superconductivity-induced changes in perpendicular magnetic anisotropy of V/MgO/Fe/MgO/Fe/Co junctions. Precise control on the of OOP or PMA in thin ferromagnetic films allowed to further reduce the size of magnetic bits when they were already reaching a critical size perpendicular; Dieny2017, although fine tuning it also became a difficult challenge Chuang2019; Yi2021; Sun2017. The main strategy to achieve an OOP magnetic orientation has usually been varying the F layer thickness and adding oxide barriers, but there is also another alternative: temperature or electromagnetic pulses (typicallly in the microwave frequency range) can decrease the energy barrier between the IP and OOP configurations, and then an appropiate magnetic field can trigger the desired reorientation Challener2009; Zhu2008; Martinez2018.
@@
In the case of our S/F junctions, all the necessary ingredients are in place to experimentally achieve the prediction of the modification of the MCA below $T_{C}$ for the OOP case Johnsen2019: a robust MCA (with cubic shape in our case, as was used in the theoretical model) and critical fields for both the IP and OOP reorientations, interfacial SOC that also induces perpendicular magnetic anisotropy Martinez2018; Yang2011, and a final but very desirable feature: we had several different samples with varying lateral sizes, which (on top of reproducibility) allows for different relative weights of the shape (IP) and surface (OOP) magnetic anisotropies. In the case of the smallest junctions shown Figure 10, there is a strong competition between the two anisotropies, which enabled the observation of the full OOP reorientation predicted in Ref.Johnsen2019. For larger junctions where the IP component dominates, we still observed weaker but robust changes suggesting a higher relative strength of the OOP anisotropy.
@@
Figure 10: (a) OOP reorientation in a $10\times 10$ $\mu\text{m}^{2}$ S/F/F junction, for different temperatures around $T_{C}$. (b) The same type of experiment for a larger ($30\times 30$ $\mu\text{m}^{2}$) junction, where an incomplete reorientation can be inferred from the TMR. (c) $H_{\text{OOP}}$ vs $T$ in different lateral sized junctions. A strong reduction of $H_{\text{OOP}}$ takes place below $T_{c}$ for the junction of size $10\times 10$ $\mu\text{m}^{2}$. (d) $\Phi_{\text{F}}$ vs $T$ at $H=0$ for four different sized junction (same legend that in the previous panel). The inset compares the angle at $H=0$ and $T=5$ K (above $T_{C}$) and at $T=0.3$ K (below $T_{C}$) for the different lateral sizes. The colored dashed lines are guides for the eyes, while the vertical, black, dashed lines mark $T_{C}$. Figure adapted from ref. GonzalezRuano2021 with the necessary permissions.
@@
In Figure 10, it can be seen how the OOP reorientation changes under different temperatures for a $10\times 10$ $\mu\text{m}^{2}$ junction. Below $T_{C}$, there is a decreasing in the applied field that triggers a full OOP magnetization reorientation ($H_{\text{OOP}}$). In junctions with larger lateral sizes, the same type of reorientation was also present below $T_{C}$, although in this case it was smaller (Figure 10b,c). This is consistent with an incomplete OOP reorientation or only some magnetic domains reorienting. Similar to the IP case, the angle between the two F layers is shown in Figure 10d.
@@
It is worth noting that the reason for the OOP TMR experiments being “assymetric” was an already known phenomenon Martinez2018, explained by a different density of lattice dislocations between the top and bottom parts of the Fe layer nearest to the V one. This behavior is inherent to most epitaxial fabrication methods, and in our case results in slightly different surface anisotropies, which in turn means that the magnetization is reoriented OOP preferentially for one of the field directions compared to the other one.
@@
As in the IP case, we also carried out micromagnetic simulations (again with MuMax3), trying to understand the possible influence of defects in the magnetization (preferentially near the surfaces of the Fe layer). We modelled GonzalezRuano2021 these defects as cells with a higher saturation magnetization ($M_{S}(\text{defects})=1.25\times M_{S}(\text{Fe})$), finding out that a density of $\sim 10^{-3}\%$) does not change the non-volatility of the OOP magnetic configuration, although it does affect the reorientation field to the IP configuration $H_{\text{OOP-IP}}$. If the defect density is further increased up to $\sim 2\times 10^{-3}\%$, this transition becomes volatile, which is not what we observe experimentally. This is an indication of the low density of magnetic defects in the junctions.
@@
The theoretical model GonzalezRuano2021 is also very similar to the one used to model the changes in the IP MCA. The free energy is computed from a Bogoliubov-de Gennes (BdG) Hamiltonian, showing that triplet generation has an energy cost that triggers the OOP reorientation below $T_{C}$. The Hamiltonian also includes a preferred IP MCA with cubic symmetry and the superconducting proximity effect. Similarly to the IP case, a new local energy minimum appears for the OOP magnetization when the temperature is lowered enough. In the same way as the previous section did for the IP MCA, the theoretical model calculates the free energy of the system from a tight-binding Bogoliubov-de Gennes (BdG) Hamiltonian, to show how the efficiency of the triplet leakage and its associated energy cost affects the OOP magnetization reorientation depending on the temperature of the system.
@@
Figure 11: (a) Free energy ($F$) vs OOP angle $\Phi_{\text{F}}$, compared to the free energy in the IP configuration ($F_{\text{IP}}$), for temperatures around $T_{C}$. (b) The OOP minimum becomes deeper, explaining the observed decrease of $H_{\text{OOP}}$. (c) When the soft F layer reorients to an OOP alignment, the SOC-mediated transformation of singlet Cooper pairs into LRTs is less efficient, which makes the superconducting condensate stronger, in turn decreasing the OOP free energy as $T$ goes below $T_{C}$. Figure adapted from ref. GonzalezRuano2021 with the necessary permissions.
@@ IV.3 Conductance anomalies: disentangling SOC and spin texture contributions to LRTs formation
For a long time, the conversion of “conventional” singlet Cooper pairs to equal-spin triplet pairs in S/F interfaces has been linked to inhomogeneous magnetism or spin textures (STs) uti1999; Bergeret2001; keizer-nature-06; Khaire2010; Robinson2010. Theoretical works Gorkov2001; Bergeret2013; Bergeret2014; Jacobsen2015 showed that LRT pair formation could also be produced by interfacial spin-orbit coupling (SOC), which as we have seen was later confirmed experimentally Banerjee2018; Martinez2020; GonzalezRuano2020; GonzalezRuano2021; Jeon2020; Cai2021.
It is well known that in clean N/F and S/F hybrids, Fabry-Pérot interferences due to electron confinement can result in the so-called MacMillan-Rowell resonances (MRRs) MRR1; MRR2, originated as two Andreev and two normal reflections taking place inside the F electrode interfacing the S deJong1995; Visani2012; Melnikov2012; Costa2022 (Figure 12a). Tomasch resonances Tomasch1966 (TRs) are similar phenomena occurring in the S layer that have been recently put forward to distinguish between conventional and topological superconductivity Strkalj2024. The above-gap conductance anomalies (CAs) arising from these two processes were suggested as experimental proof for LRTs formation in S/F heterostructures Visani2012. The full suppression of these CAs under IP magnetic fields pointed towards a direct link between LRT generation and spin textures Visani2015.
The epitaxial V/MgO/Fe heterostructures that we discuss here are an example of a S/F system coupled by interfacial SOC Martinez2020; GonzalezRuano2020; GonzalezRuano2021. As outlined previously, Rashba SOC at the Fe/MgO interface causes perpendicular magnetic anisotropy (PMA) in the interfacial atomic layers Dieny2017, which competes with the in-plane shape anisotropy Martinez2018. For this reason, one could expect that both SOC and STs could contribute to spin-triplet Cooper pair formation at relatively low magnetic fields.
@@
where $n=0,1,2,$... label the successive conductance oscillation peaks, $h$ is the Planck constant, $v_{F}^{\text{Fe}}$ is the Fermi velocity in the layer where the interference takes place (in our case, the soft Fe layer), and $t_{\text{Fe}}$ its thickness (10 nm in our S/F/F junctions and 45 nm in the F/S/F ones). For MRRs to occur, the gap-induced phase coherence between the initial electron and the reflected hole is maintained through the whole path back and forth the F layer (a total distance of at least 40 nm and 180 nm for the S/F/F and F/S/F junctions respectively). This is a strong hint of long-range superconducting proximity effects, as it means that superconducting correlations must survive deep into the Fe layer.
@@
The experiments leads us to two conclusions: First, the CAs are linked with superconductivity, since they disappear above $T_{C}$. Second, they are related to LRTs, which can survive high applied fields and inside the ferromagnet’s exchange field. We exclude resonant transmission of normal electrons through quantum well states Tao2015; Tao2019 as the main mechanism behind the observed CAs (see Ref.Tuero2024 discussion and its Supplemental Material for more details on this). Overall, the observation of CAs is indicative of the smoothness of the interfaces in the junctions ensured by the layer-by-layer epitaxial growth, which allows for the observation of quasiparticle interference effects.
To investigate the possible contribution of spin textures to the CAs and LRT formation, we have studied them in a field range not yet saturating the Fe layer magnetization. Figure 13a shows a $G(V)$ curve around a CA peak of a S/F/F junction at $T=0.3$ K after the background conductance is subtracted. This CA persisted for fields up to $H_{\text{IP}}=2$ T and $H_{\text{OOP}}=0.7$ T. The dependence of the CA with IP and OOP applied fields is presented in Fig. 13b. It remains approximately constant for an IP magnetic field up to 2 T, in qualitative agreement with experiments on smaller junctions. When an OOP field is applied, the CA is first enhanced with respect to the base level. The maximum is reached for $H_{\text{OOP}}\approx 0.3$ T, and further increasing the OOP field removes the CA excess (see Figure 13b).
To understand these results, we consider the role of PMA-induced spin textures at the MgO/Fe interface, sketched in Figure 13c. The CA amplitude under IP field gives a baseline for the spin-triplet generation rate via SOC, as the magnetization of the Fe layer in this situation is saturated thanks to the IP anisotropy. When a low $H_{\text{OOP}}$ is applied, the Fe atomic layers near the surface align with this field due to the PMA, increasing the angle $\phi$ with the IP direction. The presence of these spin textures close to the MgO/Fe interface could then open an additional channel for LRT generation Bergeret2005; Robinson2010; Buzdin2007, which would enhance the CA. When the OOP field is further increased, the inner layers of Fe (which were initially oriented IP) would also align with the OOP magnetic field, removing the STs (therefore reducing the angle gradient $\Delta\phi$) and letting the V/MgO/Fe system again only with the SOC contribution to LRT formation.
@@
Figure 13: (a) $\Delta G(V)$ of a CA for different IP and OOP applied fields in a $30\times 30~\mu\text{m}^{2}$ S/F/F junction over the background conductance. (b) Amplitude of the CA vs IP and OOP fields. Error bars mark the 95% confidence bounds. The lines are guides for the eye. The numbers within the graph correspond to the steps in panel (c), where we sketch the magnetization configuration and define the magnetization gradient (spin texture) $\Delta\phi$ simulated in the next panel. (1) After an initial IP saturation of 0.3 T, the magnetization is homogeneous. (2) A small OOP field starts inducing an angle gradient near the surface due to the PMA. (3) as the OOP field increases, the angle gradient reaches a maximum. (4) a large enough OOP field saturates the magnetization, returning to a reduced spin gradient. (d) Dependence of the spin gradient $\Delta\phi$ vs IP and OOP applied field from micromagnetic simulations. Adapted from Ref.Tuero2024 with the necessary permissions.

### v2

@@ IV.2.1 Superconductivity-induced change in the IP magnetic anisotropy
The most effective way of exploring the in-plane MCA landscape of the S/F/F junctions were magnetic field rotations, in which we used a field that was greater than the coercive field of the soft F layer but lower than the one of the hard F layer (see Figure 3b). The range of 70 to 120 Oe worked best in our case, depending on the sample. Using this range of fields ensured that the hard F layer remained fixed, while the soft one could freely rotate its magnetization following the applied field direction, revealing its own MCA in the process. This is shown in Figure 3b, where we can distinguish the two easy axes aligned with the [100] and [010] crystallographic directions as the soft F layer moves, setting the junction in the parallel P, perpendicular-in-plane (PIP) and AP magnetic configurations.
@@
We used the experimental resistances to approximate the exact angle between both F layers in any situation where we could find different resistance values (found both above and below the critical temperature of vanadium). We used these resistances, for known magnetic configurations, to calibrate a modified version of the Slonczewski model Slonczewski, which can accurately model the S/F/F junctions due to their large effective spin polarization Martinez2018; Parkin2004; Yuasa2004.
@@
Figure 8: Resistance vs in-plane magnetic field rotation angle for a V/MgO/Fe/MgO/Fe/Co junction, for different temperatures from above to below $T_{C}$, when the rotation is started from the P (a-c) and the AP (d-f) magnetic configurations, with the corresponding crystalline axis indicated. The inset helps visualizing the experiment, defining the angles between the two F layers ($\Phi_{\text{F}}$) and the one of the applied field ($\Phi_{\text{H}}$). (b) and (e) show 10 histograms from the data in (a) and (d) respectively. (c) and (f) take the data from (b) and (e) during the second half of the magnetic field rotation, representing each histogram’s count vs temperature around the intermediate (normal state hard axis) states: $\text{AP}+45^{\circ}$ or the $\overline{1}10]$ axis, $\text{AP}+90^{\circ}=\text{PIP}$ or [010], and $\text{AP}+135^{\circ}$ or [110]. Figure adapted from Ref. GonzalezRuano2020 with the necessary permissions.
@@
The magnetization rotations were then performed at different temperatures, from above to below $T_{C}$, in several samples. In Figure 8, we show above gap conductance versus applied magnetic field angle on two different samples, with the correspondent relative magnetization angle analysis. While the resulting TMR is not affected by temperature during the first half of the rotation, the second half has noticeable changes of the relative magnetization angle below $T_{C}$. In particular, we see new free layer magnetization configurations emerge in the superconducting state. According to the calibrated model, it seems to be oriented near the hard axes of the soft F layer, 45 from the easy directions. These new configurations disappear when increasing temperature and approaching $T_{C}$.
@@
Before continuing to fit the experimental results to the theoretical model, we want to discard some extrinsic effects. In the IP situation, the main candidate is the formation of domain walls (DWs). To asses their possible role in changing the resistance of the junctions during the field rotations (which could potentially invalidate our macrospin approach with the Slonczewsky model), we used the MuMax3 mumax micromagnetic simulation software to compare the formation of DWs in the [100] and [110] directions. The results showed that these DWs, in case they are present, should preferentially pin the magnetization along the [100] direction, and form less in the [110] one (see Ref. gonzalez-prb-20). Since we observe a high spin polarization when the magnetization is saturated along the [100] axis, indicating that no DWs are present, we can be reasonably confident that no DWs should appear during the rest of the rotations.
@@
Once the extrinsic factors were taken care of, we turn to the theoretical model Johnsen2019, adapting it to our system. As mentioned, the main idea is that the invariance of the superconducting proximity effect to the angle of the F layer magnetization is broken in the presence of SOC Jacobsen2015; Johnsen2019. Formally, the model is based on a tight-binding Bogolioubov-de Gennes Hamiltonian, which includes terms for electron hopping both within and between the different layers, interfacial Rashba SOC, spin exchange splitting for the Fe layers, and conventional superconductivity in vanadium. The total free energy is calculated by taking into account the magnetization-dependent superconducting proximity effect, bi-axial in-plane MCA, and a weak antiferromagnetic coupling between the F layers that was observed experimentally. The result is shown in Figure 9, where the free energy is plotted vs the angle of the magnetization ($\Phi_{\text{F}}$) under different temperatures. This nicely reproduces the new 45 easy axis below $T_{C}$. For more calculation details, seeGonzalezRuano2020.
@@
Figure 9: (a) Free energy $F$ vs in-plane magnetization angle $\Phi_{\text{F}}$ for different temperatures around $T_{C}$. (b) Sketch of the physical mechanism behind the S-induced change in magnetocrystalline anisotropy (MCA): Above $T_{C}$, V is a normal metal, while the Fe layer has a 4-fold in-plane MCA (yellow line). Below $T_{C}$, V is superconducting and influences the Fe with a magnetization-dependent penetration of Cooper pairs. The generation of triplets is lower for the $\overline{1}\overline{1}0]$ direction, maximizing the superconducting condensate energy gain. This in turn modifies the effective MCA of the Fe layer (cyan line), allowing the magnetization to reorient towards the $\overline{1}\overline{1}0]$ axis (blue arrow). Figure adapted from ref. GonzalezRuano2020 with the necessary permissions.
@@ IV.2.2 Superconductivity-assisted change of the OOP magnetic anisotropy
We turn now to describe superconductivity-induced changes in perpendicular magnetic anisotropy of V/MgO/Fe/MgO/Fe/Co junctions. Precise control over the OOP or PMA in thin ferromagnetic films allowed to further reduce the size of magnetic bits when they were already reaching a critical size perpendicular; Dieny2017, although fine tuning it also became a difficult challenge Chuang2019; Yi2021; Sun2017. The main strategy to achieve an OOP magnetic orientation has usually been varying the F layer thickness and adding oxide barriers, but there is also another alternative: temperature or electromagnetic pulses (typicallly in the microwave frequency range) can decrease the energy barrier between the IP and OOP configurations, and then an appropiate magnetic field can trigger the desired reorientation Challener2009; Zhu2008; Martinez2018.
@@
In the case of our S/F junctions, all the necessary ingredients are in place to experimentally achieve the prediction of the modification of the MCA below $T_{C}$ for the OOP case Johnsen2019. Furthermore, we have studied several different samples with varying lateral sizes, which (on top of reproducibility) allows for different relative weights of the shape (IP) and surface (OOP) magnetic anisotropies. In the case of the smallest junctions shown in Figure 10, there is a strong competition between the two anisotropies, which enabled the observation of the full OOP reorientation predicted in Ref.Johnsen2019. Even for larger junctions where the IP component dominates, we still observed weaker but robust changes suggesting a higher relative strength of the OOP anisotropy.
@@
Figure 10: (a) Out-of-plane (OOP) reorientation in a $10\times 10$ $\mu\text{m}^{2}$ S/F/F junction, for different temperatures around $T_{C}$. (b) The same type of experiment for a larger ($30\times 30$ $\mu\text{m}^{2}$) junction, where an incomplete reorientation can be inferred from the TMR. (c) $H_{\text{OOP}}$ vs $T$ in different lateral sized junctions. A strong reduction of $H_{\text{OOP}}$ takes place below $T_{c}$ for the junction of size $10\times 10$ $\mu\text{m}^{2}$. (d) $\Phi_{\text{F}}$ vs $T$ at $H=0$ for four different sized junction (same legend that in the previous panel). The inset compares the angle at $H=0$ and $T=5$ K (above $T_{C}$) and at $T=0.3$ K (below $T_{C}$) for the different lateral sizes. The colored dashed lines are guides for the eyes, while the vertical, black, dashed lines mark $T_{C}$. Figure adapted from ref. GonzalezRuano2021 with the necessary permissions.
@@
In Figure 10, it can be seen how the OOP reorientation changes under different temperatures for a $10\times 10$ $\mu\text{m}^{2}$ junction. Below $T_{C}$, there is a decreasing in the applied field that triggers a full OOP magnetization reorientation ($H_{\text{OOP}}$). In junctions with larger lateral dimensions, the same type of reorientation was observed below $T_{C}$, althoughthe effect was weaker (Figure 10b,c). This behavior is consistent with an incomplete OOP reorientation or the reorientation of only certain magnetic domains. As in the IP case, the angle between the two F layers is shown in Figure 10d.
@@
It is worth noting that the reason for the OOP TMR experiments being “asymmetric” was an already known phenomenon Martinez2018, explained by a different density of lattice dislocations between the top and bottom parts of the Fe layer nearest to the V one. This behavior is inherent to most epitaxial fabrication methods, and in our case results in slightly different surface anisotropies, which in turn means that the magnetization is reoriented OOP preferentially for one of the field directions compared to the other one.
@@
As in the IP case, we also carried out micromagnetic simulations using MuMax3, trying to understand the possible influence of defects in the magnetization (preferentially near the surfaces of the Fe layer). We modelled GonzalezRuano2021 these defects as cells with a higher saturation magnetization ($M_{S}(\text{defects})=1.25\times M_{S}(\text{Fe})$), finding out that a density of $\sim 10^{-3}\%$) does not change the non-volatility of the OOP magnetic configuration, although it does affect the reorientation field to the IP configuration $H_{\text{OOP-IP}}$. If the defect density is further increased up to $\sim 2\times 10^{-3}\%$, this transition becomes volatile, which is not what we observe experimentally. This is an indication of the low density of magnetic defects in the junctions.
@@
The theoretical model GonzalezRuano2021 is also very similar to the one used to model the changes in the IP MCA. The free energy is computed from a Bogoliubov-de Gennes (BdG) Hamiltonian, showing that triplet generation has an energy cost that triggers the OOP reorientation below $T_{C}$. The Hamiltonian also includes a preferred IP MCA with cubic symmetry and the superconducting proximity effect. Similarly to the IP case, a new local energy minimum appears for the OOP magnetization when the temperature is lowered enough.
@@
Figure 11: (a) Free energy ($F$) vs out-of-plane (OOP) angle $\Phi_{\text{F}}$, compared to the free energy in the in-plane (IP) configuration ($F_{\text{IP}}$), for temperatures around $T_{C}$. (b) The OOP minimum becomes deeper, explaining the observed decrease of $H_{\text{OOP}}$. (c) When the soft F layer reorients to an OOP alignment, the spin-orbit coupling-mediated transformation of singlet Cooper pairs into long-range triplets is less efficient, which makes the superconducting condensate stronger, in turn decreasing the OOP free energy as $T$ goes below $T_{C}$. Figure adapted from ref. GonzalezRuano2021 with the necessary permissions.
@@ IV.3 Conductance anomalies: disentangling SOC and spin texture contributions to LRTs formation
For a long time, the conversion of “conventional” singlet Cooper pairs to equal-spin triplet pairs in S/F interfaces has been linked to inhomogeneous magnetism or spin textures (STs) uti1999; Bergeret2001; keizer-nature-06; Khaire2010; Robinson2010. Theoretical works Gorkov2001; Bergeret2013; Bergeret2014; Jacobsen2015 showed that LRT pair formation could also be generated by interfacial SOC, which as we have seen was later confirmed experimentally Banerjee2018; Martinez2020; GonzalezRuano2020; GonzalezRuano2021; Jeon2020; Cai2021.
It is well known that in clean N/F and S/F hybrids, Fabry-Pérot interferences due to electron confinement can result in the so-called MacMillan-Rowell resonances (MRRs) MRR1; MRR2, originated as two Andreev and two normal reflections taking place inside the F electrode interfacing the S layer deJong1995; Visani2012; Melnikov2012; Costa2022 (Figure 12a). Tomasch resonances Tomasch1966 (TRs) are similar phenomena occurring in the S layer that have been recently put forward to distinguish between conventional and topological superconductivity Strkalj2024. The above-gap conductance anomalies (CAs) arising from these two processes were suggested as experimental proof for LRTs formation in S/F heterostructures Visani2012. The full suppression of these CAs under IP magnetic fields pointed towards a direct link between LRT generation and spin textures Visani2015.
@@
where $n=0,1,2,$... label the successive conductance oscillation peaks, $h$ is the Planck constant, $v_{F}^{\text{Fe}}$ is the Fermi velocity in the layer where the interference takes place (in our case, the soft Fe layer), and $t_{\text{Fe}}$ its thickness (10 nm in our S/F/F junctions and 45 nm in the F/S/F ones). For MRRs to occur, the gap-induced phase coherence between the initial electron and the reflected hole is maintained through the whole path back and forth in the F layer (a total distance of at least 40 nm and 180 nm for the S/F/F and F/S/F junctions, respectively). This is a strong hint of long-range superconducting proximity effects, as it means that superconducting correlations must survive deep into the Fe layer.
@@
The experiments leads us to two conclusions. First, the CAs are linked with superconductivity, since they disappear above $T_{C}$. Second, they are related to LRTs, which can survive high applied fields and inside the ferromagnet’s exchange field. We exclude resonant transmission of normal electrons through quantum well states Tao2015; Tao2019 as the main mechanism behind the observed CAs (see Ref.Tuero2024 discussion and its Supplemental Material for more details on this). Overall, the observation of CAs is indicative of the smoothness of the interfaces in the junctions ensured by the layer-by-layer epitaxial growth, which allows for the observation of quasiparticle interference effects.
To investigate the possible contribution of spin textures to the CAs and LRT formation, we have studied them in a field range not yet saturating the Fe layer magnetization. Figure 13a shows a $G(V)$ curve around a CA peak of a S/F/F junction at $T=0.3$ K after the background conductance is subtracted. This CA persisted for fields up to $H_{\text{IP}}=2$ T and $H_{\text{OOP}}=0.7$ T. The dependence of the CA with IP and OOP applied fields is presented in Figure 13b. It remains approximately constant for an IP magnetic field up to 2 T, in qualitative agreement with experiments on smaller junctions. When an OOP field is applied, the CA is first enhanced with respect to the base level. The maximum is reached for $H_{\text{OOP}}\approx 0.3$ T, and further increasing the OOP field removes the CA excess (see Figure 13b).
To understand these results, we consider the role of PMA-induced spin textures at the MgO/Fe interface, sketched in Figure 13c. The CA amplitude under IP field gives a baseline for the spin-triplet generation rate via SOC, as the magnetization of the Fe layer in this situation is saturated thanks to the IP anisotropy. When a low $H_{\text{OOP}}$ is applied, the Fe atomic layers near the surface align with this field due to the PMA, increasing the angle $\phi$ with the IP direction. The presence of these spin textures close to the MgO/Fe interface could then open an additional channel for LRT generation Bergeret2005; Robinson2010; Buzdin2007, which would enhance the CA. When the OOP field is further increased, the inner layers of Fe (which were initially oriented IP) would also align with the OOP magnetic field, removing the STs (therefore reducing the angle gradient $\Delta\phi$) and letting the V/MgO/Fe system only with the SOC contribution to LRT formation.
@@
Figure 13: (a) $\Delta G(V)$ of a conductance anomaly (CA) for different in-plane (IP) and out-of-plane (OOP) applied fields in a $30\times 30~\mu\text{m}^{2}$ S/F/F junction over the background conductance. (b) Amplitude of the CA vs IP and OOP fields. Error bars mark the 95% confidence bounds. The lines are guides for the eye. The numbers within the graph correspond to the steps in panel (c), where we sketch the magnetization configuration and define the magnetization gradient (spin texture) $\Delta\phi$ simulated in the next panel. (1) After an initial IP saturation of 0.3 T, the magnetization is homogeneous. (2) A small OOP field starts inducing an angle gradient near the surface due to the perpendicular magnetic anisotropy. (3) as the OOP field increases, the angle gradient reaches a maximum. (4) a large enough OOP field saturates the magnetization, returning to a reduced spin gradient. (d) Dependence of the spin gradient $\Delta\phi$ vs IP and OOP applied field from micromagnetic simulations. Adapted from Ref.Tuero2024 with the necessary permissions.